How to convert Time to fractions such as 7 Hrs. 20 mins have been given by 22/3 Please Help me ... I am bit confused .

Grade:11

1 Answers

Ritwik Garg
11 years ago

Its simple, convert everything to minutes and then divide by 60 to get the hours in fractions

i.e. 7 hrs 20 min = 440 min

time= 440/60 =22/3 hrs.

 

Please rate my answer.